### v1

@@ 1.1 Our results
For a fixed prime $p$, there is a non-adaptive $\mathsf{poly}(s,1/\epsilon)$ query algorithm with confidence $2/3$, which tests whether a given function $f:\mathbb{Z}_{p}^{n}\rightarrow\{-1,+1\}$, is $s$-sparse or $\epsilon$-far from $s$-sparse-Boolean.
We can also obtain a similar testing algorithm for sparsity, for a Boolean-valued function over $\mathcal{G}$. The generalized algorithm is discussed in Appendix G.
We complement our result by showing a query-complexity lower bound for sparsity-testing algorithms. Gopalan et al. [15] gave a $\Omega(\sqrt{s})$ lower bound for $s$-sparsity testing algorithms over $\mathbb{Z}_{2}^{n}$. An important component of their proof was to cleverly use an alternative notion of degree (borrowed from [7]) of a Boolean function over $\mathbb{Z}_{2}$. We also give a similar lower bound over $\mathbb{Z}_{p}^{n}$, by appropriately generalizing the useful notion of degree. For details on the definition of degree, see proof idea of Theorem 1.9 in Section 1.2 and Appendix H.
@@
For Boolean valued functions on $\mathbb{Z}_{p}^{n}$, to adaptively test $s$-sparsity, the query lower bound of any algorithm is $\Omega(\sqrt{s})$.
@@ 1.2 Proof ideas
In this section, we briefly outline the proof ideas of our main theorems. Although some of the proofs are indeed inspired by [15] (which worked over $\mathbb{Z}_{2}$), the proof techniques does not directly generalize over $\mathbb{Z}_{p}$. So, we will first discuss the proof ideas, and then clarify the differences between [15] and our techniques (see Section 1.2.1). While doing so, we will try to convey the hurdles for generalizing over $\mathbb{Z}_{p}$. Let us first sketch the proof of Theorem 1.3.
@@
    The sum of the absolute values of its coefficients is at most $p^{k}$, where $k:=\lceil\log_{p}s\rceil+1$.
@@
    The distance between the absolute value of each non-zero Fourier coefficient of $f$ and $|g(\omega_{p})|/p^{k}$ is at most $\mu/\sqrt{s}$.
To show the above, we first utilize a probabilistic method to prove that for each character $\chi_{i}\in B$ in the Fourier support of $f$, there exists a matrix $A\in\mathbb{Z}_{p}^{k\times n}$, and a column vector $b\in\mathbb{Z}_{p}^{n\times 1}$, such that – (1) $\chi_{i}$ is a solution of the linear system $A\chi=b$, and (2) no other character in the Fourier support of $f$ is a solution of $A\chi=b$, where $B$ be the set of $s$-largest Fourier coefficients of $f$. After establishing the existence of such $A$ and $b$, we consider the Fourier transform of the projection operator for the solution space of $A\chi=b$ (see Definition B.1 and Lemma B.2). The projection operator, as the name suggests, is an operator that projects $\mathbb{Z}_{p}^{n}$ onto a linear subspace which yields a partition of the Fourier spectrum of $f$. Then we show that the $\ell_{2}$ Fourier weight of $S\cap H$, i.e., $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$ is upper bounded by $\mu/\sqrt{s}$, where $S=\overline{B}$, and $H$ is a coset of $A^{\perp}$ that are solutions to the system of linear equations $A\chi=b$. For details, see Appendix B.
[…]
As stated earlier, we use a non-trivial result from algebraic number theory (Theorem 2.1), which states that if $f\in\mathbb{Z}[x]$, such that $f(\omega_{n})\neq 0$, where $\omega_{n}$ be a primitive root of unity, then, $|\prod_{i\in\mathbb{Z}_{n}^{\ast}}f(\omega_{n}^{i})|\geq 1$. We also use the fact that the sum of the absolute values of the coefficients of $g$ is at most $p^{k}$, to get an upper bound on the quantities $|g(\omega_{p}^{i})|$, for any $i\in[p-1]$. Combining these two facts, we obtain our lower bound; for details see Section B.
Proof idea of Theorem 1.4. We first show that the given function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, that is $\mu$-close to some $s$-sparse complex-valued function in $\ell_{2}$, can be written as the sum of two functions $F$ and $G$, where the Fourier coefficients of $F$ are $\lceil\log_{p}s\rceil+1$-granular and the absolute value of the Fourier coefficients of $G$ are upper bounded by $\mu/\sqrt{s}$, where $p$ is an odd prime. This follows from Theorem 1.3. Then we show that the range of $F$ is $\{-1,+1\}$, which uses the following facts:
@@
    $F^{2}$ is $12\lceil\log_{p}s\rceil+$ 1-granular.
@@
We compute $\mathbb{E}[G(x)^{2}]$ in order to find an upper bound on the Fourier coefficients of $H:=G(2f-G)$, which helps us to conclude that $\widehat{F^{2}}(\chi)=0$ for all $\chi\neq\chi_{0}$, and $\widehat{F^{2}}(\chi_{0})=1$, where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathbb{Z}_{p}^{n}$. Then we complete the proof by showing that the $\Pr_{x}[x\in\mathbb{Z}_{p}^{n}|f(x)\neq g(x)]$ is $\leq\mu^{2}/2$, which implies that $F$ is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$ by Lemma A.15. This idea has also been employed in  [15]. We have also extended this proof to a more general Abelian group $\mathcal{G}$, (see Appendix F).
Proof idea of Theorem 1.6. The example that we construct to prove Theorem 1.6 is a simple one. The crucial observation in this regard is that there are functions from $\mathbb{Z}_{p}$ to $\{-1,+1\}$, whose Fourier coefficients are smaller than $1/p$. In this work, we work with one such function $\mathbb{I}_{\geq\frac{p+1}{2}}:\mathbb{Z}_{p}\to\{-1,+1\}$. We claim that the composition function $AND_{n}\circ\mathbb{I}_{\geq\frac{p+1}{2}}$ is one such desired function from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$, such that its minimum Fourier coefficient is less than $1/p^{n}$. Here, we assume a trivial bound of $p^{n}$ on the Fourier sparsity of the function $AND_{n}\circ\mathbb{I}_{\geq\frac{p+1}{2}}$.
@@
Proof idea of Theorem 1.8. We sketch the algorithmic idea over $\mathbb{Z}_{p}^{n}$, where $p$ is an odd prime; this idea can be canonically extended to more general Abelian groups $\mathcal{G}$. The main idea of the algorithm is to partition the set of characters into buckets and estimate the weights of the individual buckets (that is the sum of the squares of the absolute values of the Fourier coefficients corresponding to the characters in the buckets). We know from Theorem 1.5 that all the coefficients of an $s$-sparse function are at least as large as $1/(p^{2}s)^{\lceil(p-1)/2\rceil}$. So, we are certain that if the weights of the buckets can be approximated within an additive error or $\tau/3$, where $\tau\geq 1/(p^{2}s)^{p-1}$, then in the case the function is $s$-sparse, not more than $s$ of the buckets can have weight more than $\tau/2$. On the other hand, we will show that if the function $f$ is $\epsilon$-far from any $s$-sparse Boolean function then with a high probability at least $(s+1)$ buckets will have weight more than $\tau$, making the estimated weight at least $32\tau/$ 3; see Lemma E.5.
@@
The challenge is that estimating the weights of the buckets is not easy if the characters the randomly partitioned into buckets. Here we use the ideas from Gopalan et al [15] and appropriately modify them to handle the technicalities of working with $\mathbb{Z}^{n}_{p}$. We choose the buckets carefully. The buckets corresponds to the cosets of $H^{\perp}$ in $\mathbb{Z}_{p}^{n}$, where $H$ is a random subspace of $\mathbb{Z}_{p}^{n}$ of dimension, $t=\Theta(s^{2})$. For such kinds of buckets, we show that estimation of the weight can be done using a small number of samples. We also need to use the concept of ‘random shift’, see Definition E.12, to avoid the corner case of characters being put into the trivial bucket.
@@
Unlike [15], it becomes a bit more challenging to prove that if $f$ is $\epsilon$-far from any $s$-sparse Boolean function over $\mathbb{Z}_{p}^{n}$, then with high probability at least $(s+1)$ buckets will have weight more than $\tau$. Since we partition the set of characters by cosets, the events whether two characters land in the same bucket (that is same coset) are not independent – since two characters (in the case $p\neq 2$) can be scalar multiple of each other; this is where some additional care is required (which was not the case in [15]). Under random shifts and case-by-case analysis, we can show that the two events are not correlated, i.e., the covariance of the corresponding indicator variables of the two events is 00; see Lemma E.13. Thus, we can use Chebyshev’s inequality and then Markov’s inequality to bound the number of buckets that can be of low weight, or in other words prove that the number of “heavy” weight buckets is more than $s+1$. Due to lack of space, the proof of Theorem 1.8 is given in Appendix E.
@@
Proof idea of Theorem 1.9. In [15], Gopalan et al. proved a query lower bound over $\mathbb{Z}_{2}^{n}$, by using a natural notion of degree of a Booelan function, denoted $\deg_{2}$. They crucially used the fact that for a Boolean function $f$ over $\mathbb{Z}_{2}$, $2^{\dim(f)}\geq s_{f}\geq 2^{\deg_{2}(f)}$; this was originally proved in [7]. To define the degree, let us consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a subspace of $\mathbb{Z}_{p}^{n}$ as defined as
@@
$V_{b,r_{1},\ldots,r_{k}}\;:=\;\{x\in\mathbb{Z}_{p}^{n}:r_{j}\cdot x\,=\,b_{j}\leavevmode\nobreak\ \pmod{p}\ \forall j\in[t]\}\;.$
Then the degree over $\mathbb{Z}_{p}$ of $f$, denoted by $\deg_{p}$, is defined in the following way.
$\deg_{p}(f)\;=\;\max_{\ell}\{\ell\;=\dim(V_{b,r_{1},\ldots,r_{t}})\;:\;s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}\;=\;p^{\dim(V_{b,r_{1},\ldots,r_{t}})}\},$
@@
where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$. [15] defined the degree over $\mathbb{Z}_{2}^{n}$ via similar restrictions. However, [15] argued that this is a natural definition of the degree over $\mathbb{Z}_{2}^{n}$. To argue, observe that $\widehat{f}=\frac{1}{2^{n}}H_{n}f$, where $H_{n}$ is the $2^{n}\times 2^{n}$ Hadamard matrix, when $x_{i}\in\mathbb{Z}_{2}^{n}$ are seen in lexicographic order. Consider the restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ that takes the value $1$ at those points such that each entry of $\widehat{f}|_{V_{b,r_{1},\ldots,r_{t}}}$ is nonzero. Then $\deg_{2}$ can be defined as the dimension of the subspace $V_{b,r_{1},\ldots,r_{t}}$ which is largest amongst them! In that case, all the Fourier coefficients of $f|_{V_{b,r_{1},\ldots,r_{t}}}$ are nonzero.
@@
Interestingly, we can also show that the above definition of $\deg_{p}$ is natural, mainly because $\widehat{f}=\frac{1}{p^{n}}V_{n}f$, where $V_{1}$ is a $p\times p$ Vandermonde matrix $V_{1}$, whose $(i,j)$-th entry is $(V_{1})_{i,j}:=\omega_{p}^{(i-1)(j-1)}$, and $V_{n}$ is defined by taking $n$ many Kronecker products of $V_{1}$, i.e., $V_{n}:=V_{1}^{\otimes n}=V_{1}\otimes\cdots\otimes V_{1}$. Note that $H_{n}=V_{n}$, over $\mathbb{Z}_{2}^{n}$. Similar to [7], one can also show that $p^{\dim(f)}\geq s_{f}\geq p^{\deg_{p}(f)}$ (Lemma I.3). This plays a crucial role in the proof.
@@
We first define two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ on the set of Boolean valued functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$. Let us choose a random $t$-dimensional subspace $H$ of $\mathbb{Z}_{p}^{Ct}$, for some parameter $C$ (to be fixed later). Let $\mathcal{C}$ be the set of all cosets of $H$. There are 2 main steps as follows.
@@
    We choose a random function $f$ randomly from $\mathbb{Z}_{p}^{Ct}$, conditioned on the fact that $f$ is $2-\tau$ far in $\ell_{2}$ from any function which has $\deg_{p}=t$, where $\tau$ is as defined in Theorem 1.8. Let us call this distribution $\mathcal{D}_{No}$.

### v2

@@ 1.1 Our results
There is a non-adaptive $\mathsf{poly}(s,1/\epsilon)$ query algorithm with confidence $2/3$, which tests whether a given function $f:\mathcal{G}\rightarrow\{-1,+1\}$, is $s$-sparse or $\epsilon$-far from $s$-sparse-Boolean.
We complement our result by showing a query-complexity lower bound for sparsity-testing algorithms. Gopalan et al. [GOS+11] gave a $\Omega(\sqrt{s})$ lower bound for $s$-sparsity testing algorithms over $\mathbb{Z}_{2}^{n}$. An important component of their proof was to cleverly use an alternative notion of degree (borrowed from [BC99]) of a Boolean function over $\mathbb{Z}_{2}$. We also give a similar lower bound over $\mathcal{G}$, by appropriately generalizing the useful notion of degree. For details on the definition of degree, see proof idea of Theorem 1.9 in Section 1.2 and in Section 7, Section 8.
@@
For Boolean valued functions on $\mathcal{G}$, to adaptively test $s$-sparsity, the query lower bound of any algorithm is $\Omega(\sqrt{s})$.
@@ 1.2 Proof ideas
In this section, we briefly outline the proof ideas of our main theorems. Note that $\mathbb{Z}_{p}^{n}$ is a vector space for prime pp. For general Abelian groups it can be shown that vector space structure does not exists. To deal with this issue we define an operation $\mathcal{G}\times\mathcal{G}\to\mathbb{Z}_{\Delta}$, which is a generalization of the inner product when $\mathcal{G}=\mathbb{Z}_{p}^{n}$. Note, that the operation, ‘*’, cannot be an inner-product in general as $\mathcal{G}$ does not have a vector space structure. We call it the “pseudo inner-product” operation (see Definition 2.12), and using this pseudo inner-product operation we define certain normal subgroups that acts like subspaces of $\mathcal{G}$, and their cosets are used to partition $\mathcal{G}$. Details are given in Section 2. Let us first sketch the proof of Theorem 1.3.
@@
    The sum of the absolute values of its coefficients is at most $\Delta^{k}$, where $k:=\lceil\log_{\Delta}2s\rceil$.
@@
    The distance between the absolute value of each non-zero Fourier coefficient of $f$ and $|g(\omega_{\Delta})|/\Delta^{k}$ is at most $\mu/\sqrt{s}$.
To show the above, we first utilize a probabilistic method to prove that for each character $\chi_{i}\in B$ in the Fourier support of $f$, there exists a matrix $A\in\mathbb{Z}_{\Delta}^{k\times T}$, and a column vector $b\in\mathbb{Z}_{\Delta}^{T\times 1}$, such that – (1) $\chi_{i}$ is a solution of the system of equations $A\chi=b$, and (2) no other character in the Fourier support of $f$ is a solution of $A\chi=b$, where $B$ be the set of $s$-largest Fourier coefficients of $f$. Using our pseudo inner product structure, we establish the existence of such AA and $b$, we consider the Fourier transform of the projection (Definition 3.1) for the solution space of $A\chi=b$ (see Definition 3.1 and Lemma 3.2). The projection operator, as the name suggests, is an operator that projects $\mathcal{G}$ onto a coset which yields a partition of the Fourier spectrum of $f$. Then we show that the $\ell_{2}$ Fourier weight of $S\cap H$, i.e., $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$ is upper bounded by $\mu/\sqrt{s}$, where $S=\overline{B}$, and $H$ is a coset of $A^{\perp}$ that are solutions to the system of linear equations $A\chi=b$. For details, see Section 3.
Proof idea of Theorem 1.5. If we put $\mu=0$ in Theorem 1.3, we get that there exists a $g\in\mathbb{Z}[X]$, with the sum of the absolute values of its coefficients is at most $\Delta^{k}$, such that $|\widehat{f}(\chi)|\geq|g(\omega_{\Delta})/\Delta^{k}|$, where $k=\lceil\log_{\Delta}2s_{f}\rceil$, $s_{f}$ being the sparsity of the Boolean valued function $f$. The remaining part of the proof is to show that for any polynomial $g$ with the aforementioned properties, $|g(\omega_{\Delta})|/\Delta^{k}\,\geq\,1/(2s_{f})^{\lceil\varphi(\Delta)/2\rceil}$.
As stated earlier, we use a non-trivial result from algebraic number theory (Theorem 3.8), which states that if $f\in\mathbb{Z}[x]$, such that $f(\omega_{n})\neq 0$, where $\omega_{n}$ be a primitive root of unity, then, $|\prod_{i\in\mathbb{Z}_{n}^{\ast}}f(\omega_{n}^{i})|\geq 1$. We also use the fact that the sum of the absolute values of the coefficients of $g$ is at most $\Delta^{k}$, to get an upper bound on the quantities $|g(\omega_{\Delta}^{i})|$, for any $i\in\{a\in\mathbb{Z}_{\Delta}:\gcd(i,a)=1\}$. Combining these two facts, we obtain our lower bound; for details see Section 3.
Proof idea of Theorem 1.4. We first show that the given function $f:G\to\{-1,+1\}$, that is $\mu$-close to some $s$-sparse complex-valued function in $\ell_{2}$, can be written as the sum of two functions $F$ and $G$, where the Fourier coefficients of $F$ are $\lceil\log_{\Delta}2s\rceil$-granular and the absolute value of the Fourier coefficients of $G$ are upper bounded by $\mu/\sqrt{s}$. This follows from Theorem 1.3. Then we show that the range of $F$ is $\{-1,+1\}$, which uses the following facts:
@@
    $F^{2}$ is $2\lceil\log_{\Delta}2s\rceil$-granular.
@@
We compute $\mathbb{E}[G(x)^{2}]$ in order to find an upper bound on the Fourier coefficients of H:=G(2fG)H:=G(2f-G), which helps us to conclude that $\widehat{F^{2}}(\chi)=0$ for all $\chi\neq\chi_{0}$, and $\widehat{F^{2}}(\chi_{0})=1$, where $\chi_{0}$ is the character which takes the value 11 at all points in $\mathcal{G}$. Then we complete the proof by showing that the $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$ is $\leq\mu^{2}/2$, which implies that $F$ is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$ by Lemma 2.43. This idea has also been employed in  [GOS+11].
@@
Proof idea of Theorem 1.8. The main idea of the algorithm is to partition the set of characters into buckets (or cosets) using our pseudo inner product structure, and estimate the weights of the individual buckets (that is the sum of the squares of the absolute values of the Fourier coefficients corresponding to the characters in the buckets). We know from Theorem 1.5 that all the coefficients of an $s$-sparse function are at least as large as $1/(2s)^{\lceil\varphi(\Delta)/2\rceil}$. So, we are certain that if the weights of the buckets can be approximated within an additive error or $\tau/3$, where $\tau\geq 1/(2s)^{\varphi(\Delta)}$, then in the case the function is $s$-sparse, not more than $s$ of the buckets can have weight more than $\tau/2$. On the other hand, we will show that if the function $f$ is $\epsilon$-far from any $s$-sparse Boolean function then with a high probability at least $(s+1)$ buckets will have weight more than $\tau$, making the estimated weight at least $2\tau/3$; see Lemma 6.5.
@@
The challenge is that estimating the weights of the buckets is not easy if the characters the randomly partitioned into buckets. Here we use the ideas from Gopalan et al [GOS+11] and appropriately modify them to handle the technicalities of working with $\mathcal{G}$. We choose the buckets carefully. The buckets corresponds to the cosets of $H^{\perp}$ in $\mathcal{G}$, where $H$ is a random subspace of $\mathcal{G}$ of dimension, $t=\Theta(s^{2})$. For such kinds of buckets, we show that estimation of the weight can be done using a small number of samples. We also need to use the concept of ‘random shift’, see Definition 6.9, to avoid the corner case of characters being put into the trivial bucket.
@@
Unlike [GOS+11], it becomes a bit more challenging to prove that if $f$ is $\epsilon$-far from any $s$-sparse Boolean function over $\mathcal{G}$, then with high probability at least $(s+1)$ buckets will have weight more than $\tau$. Since we partition the set of characters by cosets, the events whether two characters land in the same bucket (that is same coset) are not independent – since two characters (in the case $\mathcal{G}\neq\mathbb{Z}_{2}^{n}$) can be scalar multiple of each other; this is where some additional care is required (which was not the case in [GOS+11]). Under random shifts and case-by-case analysis, we can show that the two events are not correlated, i.e., the covariance of the corresponding indicator variables of the two events is 0; see Lemma 2.36. Thus, we can use Chebyshev’s inequality and then Markov’s inequality to bound the number of buckets that can be of low weight, or in other words prove that the number of “heavy” weight buckets is more than $s+1$.
@@
Proof idea of Theorem 1.9. In [GOS+11], Gopalan et al. proved a query lower bound over $\mathbb{Z}_{2}^{n}$, by using a natural notion of degree of a Boolean function, denoted $\deg_{2}$. They crucially used the fact that for a Boolean function $f$ over $\mathbb{Z}_{2}$, $2^{\dim(f)}\geq s_{f}\geq 2^{\deg_{2}(f)}$; this was originally proved in [BC99]. To define the degree, let us consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a coset of $V_{0,r_{1},\ldots,r_{t}}$ in $\mathcal{G}$ as defined as
@@
$V_{b,r_{1},\ldots,r_{k}}\;:=\;\{x\in\mathcal{G}:r_{j}\cdot x\,=\,b_{j}\leavevmode\nobreak\ \pmod{\Delta}\ \forall j\in[t]\}\;.$
Then the degree over $\mathcal{G}$ of $f$, denoted by $\deg_{\mathcal{G}}$, is defined in the following way.
$\deg_{\mathcal{G}}(f)\;=\;\max_{\ell}\{\ell\;=\dim(V_{b,r_{1},\ldots,r_{t}})\;:\;s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}\;=\;\Delta^{\dim(V_{b,r_{1},\ldots,r_{t}})}\},$
@@
where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$. [GOS+11] defined the degree over $\mathbb{Z}_{2}^{n}$ via similar restrictions. However, [GOS+11] argued that this is a natural definition of the degree over $\mathbb{Z}_{2}^{n}$. To argue, observe that $\widehat{f}=\frac{1}{2^{n}}H_{n}f$, where $H_{n}$ is the $2^{n}\times 2^{n}$ Hadamard matrix, when $x_{i}\in\mathbb{Z}_{2}^{n}$ are seen in lexicographic order. Consider the restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ that takes the value 11 at those points such that each entry of $\widehat{f}|_{V_{b,r_{1},\ldots,r_{t}}}$ is nonzero. Then $\deg_{2}$ can be defined as the dimension of $V_{b,r_{1},\ldots,r_{t}}$ which is largest amongst them! In that case, all the Fourier coefficients of $f|_{V_{b,r_{1},\ldots,r_{t}}}$ are nonzero.
@@
Interestingly, we can also show that the above definition of $\deg_{\mathcal{G}}$ is natural, mainly because $\widehat{f}=\frac{1}{|\mathcal{G}|}V_{\mathcal{G}}f$, where $V_{\ell}$ is a $p_{\ell}\times p_{\ell}$ Vandermonde matrix $V_{\ell}$, whose $(i,j)$-th entry is $(V_{\ell})_{i,j}:=\omega_{p_{\ell}}^{(i-1)(j-1)}$, and $V_{\mathcal{G}}$ is defined by taking Kronecker products of $V_{1},\ldots,V_{T}$, i.e., $V_{\mathcal{G}}:=V_{1}\otimes\cdots\otimes V_{T}$. Note that $H_{T}=V_{T}$, over $\mathbb{Z}_{2}^{T}$. Similar to [BC99], one can also show that $\Delta^{\dim(f)}\geq s_{f}\geq\Delta^{\deg_{\mathcal{G}}(f)}$ (see Lemma 8.3). This plays a crucial role in the proof.
@@
We first define two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ on the set of Boolean valued functions from $\mathcal{G}$ to $\{-1,+1\}$. Let us choose a random $t$-dimensional coset $H$ of $\mathbb{Z}_{\Delta}^{Ct}$, for some parameter $C$ (to be fixed later). Let $\mathcal{C}$ be the set of all cosets of $H$. There are 2 main steps as follows.
@@
    We choose a random function $f$ randomly from $\mathbb{Z}_{\Delta}^{Ct}$, conditioned on the fact that $f$ is $2-\tau$ far in $\ell_{2}$ from any function which has $\deg_{\mathcal{G}}=t$, where $\tau$ is as defined in Theorem 1.8. Let us call this distribution $\mathcal{D}_{No}$.